OpenStudy (anonymous):

Missy knows that she needs $40,000 for a down payment on a house. She found an investment that earns 3.05% interest compounding monthly. She would like to purchase the home in 5 years. How much should she put in the account now to ensure she has her down payment?

OpenStudy (anonymous):

Let the amount of money Missy have to put in account now to be x in 5 years there will be 60 month by using geometric progression where the n th term = a*r^(n-1) and in this case, a=x , 60th term is 40000 and n = 60 r= 1.0305 therefore, 40000= x * (1.0305)^59 x=6795.61 So she will need to put 6795.61 into the account.